# An end-to-end generative diffusion model for heavy-ion collisions

### v1

@@ Abstract
transforms the prior distribution back into the data distribution by gradually removing the noise. Here, $\boldsymbol{w}$ and $\bar{\boldsymbol{w}}$ both represent the standard Wiener processes (Gaussian white noise), with ${\bf f}(\boldsymbol{x},t)$ the drift coefficient and $g(t)$ the diffusion coefficient of $\boldsymbol{x}(t)$. In the generative diffusion model, the probability distribution $p_{t}(\boldsymbol{x})$, hence the score function $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$, is generally unknown, which can be estimated by a neural network with parameter $\boldsymbol{\theta}$ via minimizing the explicit score matching (ESM) loss $\mathcal{L}_{t}^{\text{ESM}}\equiv\mathbb{E}_{p(\boldsymbol{x}_{t})}||\boldsymbol{s}_{\boldsymbol{\theta}}(\boldsymbol{x},t)-\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})||^{2}$. Once we have a trained $s_{\theta}(\boldsymbol{x},t)$, the trajectory from the prior distribution to the real data distribution can be determined following Eq. 2.
In DMs, the goal is to estimate the score function by minimizing $\mathcal{L}_{t}^{\text{ESM}}$. Even without the knowledge of the probability distribution, fortunately, one is allowed to use the denoising score matching loss (DSM), $\mathcal{L}_{t}^{\text{DSM}}\equiv\mathbb{E}_{p(\boldsymbol{x}_{t},\boldsymbol{x}_{0})}||\boldsymbol{s}_{\boldsymbol{\theta}}(\boldsymbol{x},t)-\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})||^{2}$, where the conditional probability distributions $p(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})$ is known. The DSM has a constant difference from the explicit one [67], which leaves the optimized parameter $\boldsymbol{\theta}$ unchanged, i.e.,
$\boldsymbol{\theta}^{*}$ $=\text{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{t}^{\text{DSM}}=\text{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{t}^{\text{ESM}}.$ (3)
In this Letter, we employ the so-called variance-preserving SDE [49, 47, 50], in which with ${\bf f}(\boldsymbol{x},t)=-\frac{1}{2}\beta(t)\boldsymbol{x}$ and $g(t)=\sqrt{\beta(t)}$ the conditional probability distribution is analytically solvable. After discretization, the forward SDE can be viewed as a Markov chain,
@@
where $t_{i}=i/N$, and $N$ the number of steps. With respect to the properties of Wiener process, $\sqrt{\Delta t}\boldsymbol{\varepsilon}(t_{i})$ follows a Gaussian noise of zero mean and width $\Delta t$, namely, $\mathcal{N}(0,\Delta t)$. Let $\beta_{i}=N\beta(t_{i}),\boldsymbol{x}(t_{i})=\boldsymbol{x}_{i},\boldsymbol{\varepsilon}(t_{i})=\boldsymbol{\varepsilon}_{i}$, one has $p(\boldsymbol{x}_{i}|\boldsymbol{x}_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{i}}\boldsymbol{x}_{0},\sigma_{i}^{2}\boldsymbol{I})$, where $\alpha_{i}=1-\beta_{i}$, $\bar{\alpha}_{i}=\prod_{j=0}^{i}\alpha_{j}$ and $\sigma_{i}=\sqrt{1-\bar{\alpha}_{i}}$. Therefore, $\boldsymbol{\theta}^{*}=\text{argmin}_{\boldsymbol{\theta}}\mathbb{E}_{p(\boldsymbol{x}_{t},\boldsymbol{x}_{0})}||\boldsymbol{s}_{\boldsymbol{\theta}}(\boldsymbol{x},t)+\frac{\boldsymbol{x}_{i}-\sqrt{\bar{\alpha}_{i}}\boldsymbol{x}_{0}}{\sigma_{i}^{2}}||^{2}=\text{argmin}_{\boldsymbol{\theta}}\mathbb{E}_{p(\boldsymbol{x}_{t},\boldsymbol{x}_{0})}||\boldsymbol{\varepsilon}_{\boldsymbol{\theta}}(\boldsymbol{x},t)-\boldsymbol{\varepsilon}||^{2},$ where $\boldsymbol{x}_{i}=\sqrt{\bar{\alpha}_{i}}\boldsymbol{x}_{0}+\sigma_{i}\boldsymbol{\varepsilon}$ and $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,\boldsymbol{I})$. For simplicity, we have defined the scaled score function $\boldsymbol{\varepsilon}_{\boldsymbol{\theta}}=-\sigma_{i}\boldsymbol{s}_{\boldsymbol{\theta}}$ as the optimization target, referred to as the noise prediction network.
@@
which converts the probabilistic models to the deterministic models. Additionally, fast sampling can be performed through the numerical methods of ODE [68, 69, 70].
@@
A generative diffusion model for heavy-ion collisions.— In this work, we train a generative diffusion model to function as a heavy-ion collision event generator. We carried out (2+1)D minimum bias simulations of Pb-Pb collisions at 5.02 TeV, choosing the shear viscosity $\eta/s$ to be one of three distinct values: 0.0, 0.1, and 0.2. For each value of $\eta/s$, we generate 12,000 pairs of initial entropy density profiles and final particle spectra, corresponding to 12,000 simulated events, as the training dataset. 70% of the total events are used for training and the rest are used for validation.
@@
Figure 1: The workflow of DiffHIC. The brown block is the core noise prediction network, which is a typical U-net architecture. The blue and orange boxes are the up and down sampling blocks, respectively. Both are realized by ResNet [74]. The boxes with "Att" represent the block performing the spatial attention computations [75]. The green box is the bottleneck block.
@@
Figure 2: The centrality dependence of integrated anisotropy flow. The filled symbols are the ground truth. The first column is the ideal hydrodynamic results. The second and third columns present the results with $\eta/s=0.1,\eta/s=0.2$, respectively.
Figure 3: The $p_{T}$ dependence of anisotropy flow, across all centralities. The filled symbols are the ground truth. The first row is the ideal hydrodynamic results. The second and third rows present the results with $\eta/s=0.1,\eta/s=0.2$, respectively. In each plot, the red, blue, and green lines represent $v_{2}(p_{T}),v_{3}(p_{T}),$ and $v_{4}(p_{T})$, respectively.
@@
Model Performance.— To evaluate the performance of our trained DiffHIC model, we conducted additional 10,000-event simulations for each value of $\eta/s$ across all centralities. We assess the efficacy of DiffHIC by comparing its outputs with the original numerical simulations, which were regarded as the ground truth. In terms of the generated particle spectrum, which in practice is a two-dimensional discretized distribution of pixels of size $64\times 64$ (in $p_{T}$ and $\phi_{p}$), perfect comparison between DIffHIC and the ground truth is realized with impurities appear only randomly at a level of several pixels.
@@
Comparisons involving other observables, including in particular correlations and fluctuations among these flow observables [1, 2, 76, 33, 26, 77, 78, 79], are detailed in the Supplemental Material. We first present here the integrated flow and differential flow. Owing to event-by-event fluctuations, anisotropic flow can be measured with respect to an event plane, or from multi-particle cumulants. In Fig. 2, the integrated flow of order $n=2,3,4$ are shown for all centralities, from the event-plane method (top panels), and two-particle cumulant (middle panels) and four-particle cumulant (bottom panels). Comparing to the ground truth (filled symbols), DiffHIC (colored bands) reproduces in general the predictions of the integrated flow, with slight deviation visible for $n=3$ in the four-particle cumulant. The deviation gets reduced as $\eta/s$ increases. Although differential flow signatures are normally more demanding for model characterizations, in Fig. 3, one observes a perfect comparison between the ground truth (filled symbols) and the DiffHIC predictions (colored bands) for the $p_{T}$-dependent anisotropic flow, through all centralities.
@@
To explicitly demonstrate the performance of DiffHIC, the ratios of observables, comparing the generative model to traditional simulations, are shown in Fig. 4. From left to right, the number of particle correlations involved increases, as indicated by the different color regions. These ratios are close to unity, indicating the validity of DiffHIC. Although the model precision decreases systematically as the number of correlated particles increases, it is to some extent anticipated since our simulations are limited by the spectrum resolution of $64\times 64$. A higher-resolution model would capture more details of the heavy-ion dynamics. Similar issues also exist in the high-$p_{T}$ region and peripheral collisions, where effective pixels of the spectrum reduce significantly. Another limitation is the model may struggle in other systems with different collision energies, especially for high-order fluctuations and correlations. Nonetheless, improvement should be achieved by training a model with better resolution. Given the efficiency and flexibility of DiffHIC, fine-tuning our trained model in the new dataset or retraining a specific model is also a possible solution.
@@
Conclusions and Outlook.— In this Letter, the state-of-art generative model (DiffHIC) is for the first time trained to generate the final particle spectra, from event-by-event initial entropy density profiles. While being capable of capturing the two-dimensional distribution in the momentum space accurately, as an end-to-end model, DiffHIC speeds up the numerical simulations by a factor of roughly $10^{5}$, compared to the traditional approach. Consequently, DiffHIC alleviates time and resource concerns. The DifHIC model aims to be applicable to high-precision experimental measurements, establishing a solid theoretical foundation for interpreting observables based on a huge amount of collision data, particularly in relation to the fine structures present in the systems. These include features such as nuclear deformation and the equation of state in the QGP.
@@
Acknowledgements.— Jing-An Sun thanks Xiangyu Wu, Lipei Du, and Han Gao for helpful discussions. This work is funded in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048 and SAPIN-2024-00026], in part by the China Scholarship Council, and in part by the National Natural Science Foundation of China under Grants No. 12375133 and 12147101. The dataset preparations are made on the Beluga supercomputer system at McGill University, managed by Calcul Quebec and the Digital Research Alliance of Canada. The operation of this supercomputer is funded by the Canada Foundation for Innovation (CFI), Ministère de l’Économie, des Sciences et de l’Innovation du Québec (MESI) and le Fonds de recherche du Québec - Nature et technologies (FRQ-NT).
Appendix A supplemental material
A.1 The model parameters used when preparing the datasets
Initial Stage:
We choose the most probable values from the Bayesian analysis, $k=1.0,w=1.0,p=0.0$. The norm factors are set to $18.8$ for PbPb@5.02TeV. The grid size is $[-15,15]$ and $100$
Hydrodynamic Stage:
We turn off the bulk viscosity during the evolution. The initial time is 0.4 fm. The freezeout energy density is chosen at 0.18 GeV/fm3. To view the effect of the shear viscosity, we run simulations at three different values, $0.0,0.1,0.2$. The effect of bulk viscosity is neglected.
Particlization and afterburner:
The default parameters in UrQMD are used.
A.2 The charged particle spectra
In momentum space, the azimuthal emitted particle distribution can be written as
${\frac{1}{2\pi}}\frac{d^{2}N}{p_{T}dp_{T}d\phi}=\frac{1}{2\pi}\frac{dN}{p_{T}dp_{T}}\left(1+2\sum_{n=1}^{\infty}v_{n}(p_{T})\cos[n(\phi-\Psi_{n}(p_{T}))]\right),$ (6)
where $v_{n}(p_{T})$ is the $n$-th order anisotropic flow coefficient and $\Psi_{n}(p_{T})$ is its corresponding flow plane angle. Both are $p_{T}$ differential. One can also perform Fourier expansion for the $p_{T}$ integrated spectra,
$\frac{dN}{d\phi}\propto\left(1+2\sum_{n=1}^{\infty}v_{n}\cos[n(\phi-\Psi_{n})]\right),$ (7)
where $v_{n}$ is the $n$-th order integrated anisotropic flow and $\Psi_{n}$ is its corresponding flow plane angle.
We present charged particle transverse momentum spectra $dN/2\pi p_{T}dp_{T}$ across all centralities in Fig. 5. It is shown that the model predictions (color lines) perfectly agree with the ground truth ( black circles).
Figure 5: The charged particle transverse momentum spectra. The black points are the ground truth.
Figure 6: The normalized mixed harmonic cumulants are presented, with open markers indicating the ground truth and colored bands representing the generated results. The first, second, and third columns correspond to the 4-, 6-, and 8-particle cumulants, respectively. The shear viscosity values, from left to right, are $\eta/s=0.0,0.1,$ and $0.2$.
A.3 Mixed harmonic coeffients
The mixed harmonic coefficients (MHC) have been proposed to quantify the correlation strength between different orders of flow coefficients and various moments, as they are expected to be more sensitive to medium properties and initial state correlations [80, 81, 82]. Additionally, MHC excludes symmetry plane correlations, making it less sensitive to non-flow contaminations.
The 4-particle mixed harmonic cumulants (MHC) are defined as,
$MHC(v_{n}^{2},v_{m}^{2})=\langle v_{n}^{2}v_{m}^{2}\rangle-\langle v_{n}^{2}\rangle\langle v_{m}^{2}\rangle$ (8)
The 6-particle cumulants are defined as,
$MHC(v_{2}^{4},v_{3}^{2})$ $=\langle v_{2}^{4}v_{3}^{2}\rangle-4\langle v_{2}^{2}v_{3}^{2}\rangle\langle v_{2}^{2}\rangle-\langle v_{2}^{4}\rangle\langle v_{3}^{2}\rangle+4\langle v_{2}^{2}\rangle^{2}\langle v_{3}^{2}\rangle,$
$MHC(v_{2}^{2},v_{3}^{4})$ $=\langle v_{2}^{2}v_{3}^{4}\rangle-4\langle v_{2}^{2}v_{3}^{2}\rangle\langle v_{3}^{2}\rangle-\langle v_{2}^{2}\rangle\langle v_{3}^{4}\rangle+4\langle v_{2}^{2}\rangle\langle v_{3}^{2}\rangle^{2},$
$MHC(v_{2}^{2},v_{3}^{2},v_{4}^{2})$ $=\langle v_{2}^{2}v_{3}^{2}v_{4}^{2}\rangle-\langle v_{2}^{2}v_{3}^{2}\rangle\langle v_{4}^{2}\rangle-\langle v_{2}^{2}v_{4}^{2}\rangle\langle v_{3}^{2}\rangle-\langle v_{3}^{2}v_{4}^{2}\rangle\langle v_{2}^{2}\rangle+2\langle v_{2}^{2}\rangle\langle v_{3}^{2}\rangle\langle v_{4}^{2}\rangle,$
The 8-particle cumulants are defined as,
$MHC(v_{2}^{6},v_{3}^{2})$ $=\langle v_{2}^{6}v_{3}^{2}\rangle-9\langle v_{2}^{4}v_{3}^{2}\rangle\langle v_{2}^{2}\rangle-\langle v_{2}^{6}\rangle\langle v_{3}^{2}\rangle-9\langle v_{2}^{4}\rangle\langle v_{2}^{2}v_{3}^{2}\rangle-36\langle v_{2}^{2}\rangle^{3}\langle v_{3}^{2}\rangle$
$+18\langle v_{2}^{2}\rangle\langle v_{3}^{2}\rangle\langle v_{2}^{4}\rangle+36\langle v_{2}^{2}\rangle^{2}\langle v_{2}^{2}v_{3}^{2}\rangle$
$MHC(v_{2}^{4},v_{4}^{4})$ $=\langle v_{2}^{4}v_{4}^{4}\rangle-4\langle v_{2}^{4}v_{2}^{2}\rangle\langle v_{3}^{2}\rangle-4\langle v_{2}^{2}v_{4}^{4}\rangle\langle v_{2}^{2}\rangle-\langle v_{2}^{4}\rangle\langle v_{4}^{4}\rangle-8\langle v_{2}^{2}v_{2}^{2}\rangle^{2}-24\langle v_{2}^{2}\rangle^{2}\langle v_{3}^{2}\rangle^{2}+4\langle v_{2}^{2}\rangle\langle v_{4}^{4}\rangle$
$+4\langle v_{2}^{4}\rangle\langle v_{3}^{2}\rangle^{2}+32\langle v_{2}^{2}\rangle\langle v_{3}^{2}\rangle\langle v_{2}^{2}v_{3}^{2}\rangle$
$MHC(v_{2}^{2},v_{3}^{6})$ $=\langle v_{2}^{2}v_{3}^{6}\rangle-9\langle v_{2}^{2}v_{3}^{4}\rangle\langle v_{3}^{2}\rangle-\langle v_{3}^{6}\rangle\langle v_{2}^{2}\rangle-9\langle v_{3}^{4}\rangle\langle v_{2}^{2}v_{3}^{2}\rangle-36\langle v_{2}^{2}\rangle\langle v_{3}^{2}\rangle^{3}$
$+18\langle v_{2}^{2}\rangle\langle v_{3}^{2}\rangle\langle v_{3}^{4}\rangle+36\langle v_{3}^{2}\rangle^{2}\langle v_{2}^{2}v_{3}^{2}\rangle,$
To eliminate the flow magnitude effect, the normalized cumulants can be used.
$nMHC(v_{m}^{k},v_{n}^{l})$ $=\frac{MHC(v_{m}^{k},v_{n}^{l})}{\langle v_{m}^{k}\rangle\langle v_{n}^{l}\rangle},$ (9)
$nMHC(v_{m}^{k},v_{n}^{l},v_{p}^{q})$ $=\frac{MHC(v_{m}^{k},v_{n}^{l},v_{p}^{q})}{\langle v_{m}^{k}\rangle\langle v_{n}^{l}\rangle\langle v_{p}^{q}\rangle}.$ (10)
In Fig. 6, we compare the ground truth and the generated results for various flow correlations. Except for the $(v_{2}^{2},v_{3}^{2})$ correlation at the very last peripheral bin, the agreements are excellent.
A.4 The mean transverse momentum correlations
The event-by-event mean transverse momentum fluctuations are other probes, which are sensitive to initial state fluctuations [83, 36, 84].
The second and third-order $p_{T}$ correlators are defined as,
$\langle\delta p_{T}\delta p_{T}\rangle$ $=\left\langle\frac{\sum_{i\neq j}(p_{i}-\langle\langle p_{T}\rangle\rangle)(p_{j}-\langle\langle p_{T}\rangle\rangle)}{N_{ch}(N_{ch}-1)}\right\rangle_{\text{ev}}$ (11)
$\langle\delta p_{T}\delta p_{T}\delta p_{T}\rangle$ $=\left\langle\frac{\sum_{i\neq j\neq k}(p_{i}-\langle\langle p_{T}\rangle\rangle)(p_{j}-\langle\langle p_{T}\rangle\rangle)(p_{k}-\langle\langle p_{T}\rangle\rangle)}{N_{ch}(N_{ch}-1)(N_{ch}-2)}\right\rangle_{\text{ev}}.$ (12)
In a single event, $p_{i}$ is the transverse momentum of $i-$th particle. $N_{ch}$ is the total number of charged particles. $\langle\langle p_{T}\rangle\rangle$ is the event-averaged mean transverse momentum. For efficient computations, in each event, we define
$Q_{n}=\sum_{i=1}^{N_{ch}}(p_{i})^{n}.$ (13)
The $\langle p_{T}\rangle$ and sums over pairs and triplets of particles can be expressed simply in terms of $Q_{n}$ ($n=1,2,3$):
$\langle p_{T}\rangle$ $=\frac{Q_{1}}{N_{ch}}$ (14)
$\sum_{i\neq j}p_{i}p_{j}$ $=(Q_{1})^{2}-Q_{2},$ (15)
$\sum_{i\neq j\neq k}p_{i}p_{j}p_{k}$ $=(Q_{1})^{3}-3Q_{2}Q_{1}+2Q_{3}.$ (16)
Thus, the correlators can be simplified as,
$\langle\delta p_{T}\delta p_{T}\rangle$ $=\left\langle\frac{Q_{1}^{2}-Q_{2}}{N_{ch}(N_{ch}-1)}\right\rangle-\left\langle\frac{Q_{1}}{N_{ch}}\right\rangle^{2}$ (17)
$\langle\delta p_{T}\delta p_{T}\delta p_{T}\rangle$ $=\left\langle\frac{({Q_{1}})^{3}-3Q_{2}{Q_{1}}+2Q_{3}}{N_{ch}(N_{ch}-1)(N_{ch}-2)}\right\rangle-3\left\langle\frac{({Q_{1}})^{2}-Q_{2}}{N_{ch}(N_{ch}-1)}\right>\left<\frac{{Q_{1}}}{N_{ch}}\right>+2\left<\frac{{Q_{1}}}{N_{ch}}\right\rangle^{3}.$ (18)
Fig.7 presents the centrality dependence of $\langle\delta p_{T}\delta p_{T}\rangle$ (left panel) and $\langle\delta p_{T}\delta p_{T}\delta p_{T}\rangle$ (right panel). The trained model calculations are highly consistent with the numerical simulations implying that the underlying energy density fluctuations and flow velocity fluctuations are well captured by the model.
Figure 7: The centrality dependence of 2-particle and 3-particle $p_{T}$ correlator. The open marks are the ground truth and the color bands are the generated results.
A.5 The flow and transverse momentum correlations
In the past few years, the correlations between flow and mean transverse momentum have been put forward to infer the nuclear shape [37, 38, 39, 85]. It can be calculated as,
$\rho_{n}=\frac{\langle v_{n}^{2}\{2\}\langle p_{T}\rangle\rangle-\langle v_{n}^{2}\{2\}\rangle\langle\langle p_{T}\rangle\rangle}{\sigma_{v_{n}^{2}\{2\}}\sigma_{\langle p_{T}\rangle}},$ (19)
where $\sigma_{v_{n}^{2}\{2\}},\sigma_{\langle p_{T}\rangle}$ are the standard variance of $v_{n}^{2}\{2\}$ and $\langle p_{T}\rangle$, respectively. Fig. 8 shows that the ground truth and the generated results are in good agreement except for the very last peripheral bin where hydrodynamics may not fully apply.
Figure 8: The correlation between flow $v_{2}\{2\},v_{3}\{2\},v_{4}\{2\}$ and mean transverse momentum, with centrality dependence. The open marks represent the ground truth, while the colored lines indicate the generated results. This serves as a crucial probe into nuclear structure. From left to right, the shear viscosity values are $\eta/s=0.0,0.1,$ and $0.2$, respectively.

### v2

@@ Abstract
[…]
While the goal in DMs is to estimate the score function, instead of minimizing $\mathcal{L}_{t}^{\text{ESM}}$, it is more feasible to solve from the denoising score matching loss (DSM), $\mathcal{L}_{t}^{\text{DSM}}\equiv\mathbb{E}_{p(\boldsymbol{x}_{t},\boldsymbol{x}_{0})}||\boldsymbol{s}_{\boldsymbol{\theta}}(\boldsymbol{x},t)-\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})||^{2}$, because the conditional probability $p(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})$ is solvable by construction. Noticing that ESM and DSM differs by a shift that is independent from the parameter $\boldsymbol{\theta}$}, minimization
$\boldsymbol{\theta}^{*}$ $=\text{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{t}^{\text{DSM}}=\text{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{t}^{\text{ESM}}\,$, (3)
equivalently determines the score function. The equivalence can be intuitively understood from the relation $p_{t}({\bf x})=\int d{\bf x}_{0}p({\bf x}|{\bf x}_{0})$, where the parameters do not contribute. In this Letter, we employ the so-called variance-preserving SDE [49, 47, 50], in which with ${\bf f}(\boldsymbol{x},t)=-\frac{1}{2}\beta(t)\boldsymbol{x}$ and $g(t)=\sqrt{\beta(t)}$ the conditional probability distribution is analytically solvable. After discretization, the forward SDE can be viewed as a Markov chain,
@@
[…]
@@
which converts the probabilistic models to the deterministic models. Because the reserve ODE and the SDE share the same probability distribution (cf. the Supplemential Material [67]), the sampling can be achieved identically, in terms of, for instance, fidelity, precision and full variability of the data.
Moreover, fast sampling can be performed through the numerical methods of ODE [68, 69, 70].
@@
The model generates a dataset which covers the transverse momentum $p_{T}\in[0.0,2.0]$ GeV and the rapidity $y\in[-0.5,0.5]$. These ranges are chosen to align with the kinematic regions commonly studied in heavy-ion collision experiments, ensuring that the model captures the essential features of particle production.
A generative diffusion model for heavy-ion collisions.— In this work, we train a generative diffusion model to function as a heavy-ion collision event generator. We carried out (2+1)D minimum bias simulations of Pb-Pb collisions at 5.02 TeV, choosing the shear viscosity $\eta/s$ to be one of three distinct values: 0.0, 0.1, and 0.2. For each value of $\eta/s$, we generate 12 000 pairs of initial entropy density profiles and final particle spectra, corresponding to 12 000 simulated events across all centralities, as the training dataset. The detailed parameters for the initial stage, hydrodynamic evolution, and the particlization are provided in the Supplemental Material [74]. As is a standard practice in machine learning, 70% of the total events are used for training and the rest are used for validation.
@@
Figure 1: The workflow of DiffHIC. The brown block is the core noise prediction network, which is a typical U-net architecture. The blue and orange boxes are the up and down sampling blocks, respectively. Both are realized by ResNet [75]. The boxes with "Att" represent the block performing the spatial attention computations [76]. More details of the architecture are given in the Supplemental Material [77]. The green box is the bottleneck block.
@@
Figure 2: The centrality dependence of integrated anisotropy flow of the 2nd ($v_{2}$), 3rd ($v_{3}$) and the 4th ($v_{4}$) order. The filled symbols are the ground truth. The first column is the ideal hydrodynamic results. The second and third columns present the results with $\eta/s=0.1,\eta/s=0.2$, respectively.
Figure 3: The $p_{T}$ dependence of anisotropy flow, across all centralities. The filled symbols are the ground truth. The first row is the ideal hydrodynamic results. The second and third rows present the results with $\eta/s=0.1,\eta/s=0.2$, respectively. In each plot, the red, blue, and green lines represent flow of the 2nd order ($v_{2}(p_{T}))$, the 3rd order ($v_{3}(p_{T})$) and the 4th order ($v_{4}(p_{T})$), respectively.
@@
Model Performance.— To evaluate the performance of our trained DiffHIC model, we conducted additional 10,000-event simulations for each value of $\eta/s$ in every centrality. We assess the efficacy of DiffHIC by comparing its outputs with the original numerical simulations, which were regarded as the ground truth. In terms of the generated particle spectrum, which in practice is a two-dimensional discretized distribution of pixels of size $64\times 64$ (in $p_{T}$ and $\phi_{p}$), perfect comparison between DIffHIC and the ground truth is realized with impurities appear only randomly at a level of several pixels.
@@
Comparisons involving other observables, including in particular correlations and fluctuations among these flow observables [1, 2, 78, 33, 26, 79, 80, 81], are detailed in the Supplemental Material [82]. We first present here the integrated flow and differential flow. Owing to event-by-event fluctuations, anisotropic flow can be measured with respect to an event plane, or from multi-particle cumulants. In Fig. 2, the integrated flow of order $n=2,3,4$ are shown for all centralities, from the event-plane method (top panels), and two-particle cumulant (middle panels) and four-particle cumulant (bottom panels). Comparing to the ground truth (filled symbols), DiffHIC (colored bands) reproduces in general the predictions of the integrated flow, with slight deviation visible for $n=3$ in the four-particle cumulant. The deviation gets reduced as $\eta/s$ increases. Although differential flow signatures are normally more demanding for model characterizations, in Fig. 3, one observes a perfect comparison between the ground truth (filled symbols) and the DiffHIC predictions (colored bands) for the $p_{T}$-dependent anisotropic flow, through all centralities.
@@
To explicitly demonstrate the performance of DiffHIC, the ratios of observables, comparing the generative model to traditional simulations, are shown in Fig. 4. From left to right, the number of particle correlations involved increases, as indicated by the different color regions. These ratios are close to unity, indicating the validity of DiffHIC. Although the model precision decreases systematically as the number of correlated particles increases, it is to some extent anticipated since our simulations are limited by the spectrum resolution of $64\times 64$ [^1] The choice of a $\times64$ grid resolution was driven by a careful balance between computational efficiency (GPU memory constraints) and physical fidelity (sufficient to accurately characterize particle spectra). While this resolution provides a robust foundation for our current analyses, we expect that finer resolutions could potentially yield improved precision, especially for higher-order correlations, because it captures finer details of the particle distributions. . A higher-resolution model would capture more details of the heavy-ion dynamics. Similar issues also exist in the high-$p_{T}$ region and peripheral collisions, where effective pixels of the spectrum reduce significantly. Another limitation is the model may struggle in other systems with different collision energies, especially for high-order fluctuations and correlations. Nonetheless, improvement should be achieved by training a model with better resolution. Given the efficiency and flexibility of DiffHIC, fine-tuning our trained model in the new dataset or retraining a specific model is also a possible solution.
@@
Conclusions and Outlook.— In this Letter, the state-of-art generative model (DiffHIC) is for the first time trained to generate the final particle spectra, from event-by-event initial entropy density profiles. While being capable of capturing the two-dimensional distribution in the momentum space accurately, as an end-to-end model, DiffHIC speeds up the numerical simulations by a factor of roughly $10^{5}$, compared to the traditional approach. Consequently, DiffHIC alleviates time and resource concerns.
The DifHIC model aims to be applicable to high-precision experimental measurements, establishing a solid theoretical foundation for interpreting observables based on a huge amount of collision data, and physical parameters. Notably, DiffHIC is designed to be self-adaptive, enabling the extension of its current parameter set and parameter phase space to incorporate fine structures observed in realistic systems, such as nuclear deformation [83].
@@
Acknowledgements.— Jing-An Sun thanks Xiangyu Wu, Lipei Du, and Han Gao for helpful discussions. This work is funded in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048 and SAPIN-2024-00026], in part by the China Scholarship Council, and in part by the National Natural Science Foundation of China under Grants No. 12375133 and 12347106. The dataset preparations are made on the Beluga supercomputer system at McGill University, managed by Calcul Quebec and the Digital Research Alliance of Canada. The operation of this supercomputer is funded by the Canada Foundation for Innovation (CFI), Ministère de l’Économie, des Sciences et de l’Innovation du Québec (MESI) and le Fonds de recherche du Québec - Nature et technologies (FRQ-NT).